# Contextual Status Updating: How to Maximize Situational Awareness?

### v1

@@ V-C Asymptotic Optimality
$\Psi_{\pi}((\mathbf{v}^{n})_{n=1}^{N})=$
$\mathbb{E}[(\mathbf{V}^{n}(t+1))_{n=1}^{N}|(\mathbf{V}^{n}(t))_{n=1}^{N}=(\mathbf{v}^{n})_{n=1}^{N}].$ (35)
@@ VI Discussions
To understand this characteristic, we do the following experiment: consider a safety-critical system where $N$ agents (e.g., cars) are moving in a region illustrated in Figure 3. This region is equally divided into 400 positions and the received observation $X_{n,t}$ of agent $n$ is represented by the position $x_{n}=(x_{n,1},x_{n,2})$ at time $t$. The safety level $Y_{n,t}$ is divided into three regions: {safe, cautious, dangerous}. An agent $n$ can randomly move in any of the four directions: up, down, left, and right with equal probability 0.2. If agent $n$ is in the leftmost position, then moving left means it will stay in the same position, similar criteria are applied for the rightmost, upmost, and downmost positions. The losses considered in this experiment are the same as in Example 1: $L$(dangerous, safe) $=1000,L$(safe, dangerous) $=5,L$(cautious, safe) $=10,L$(safe, cautious) $=1,1,L$(dangerous, cautious) $=100$, $L$(cautious, dangerous)$=5$, and $L$(dangerous, dangerous) $=L$(cautious, cautious) $=L$(safe, safe) $=0$.
Figure 5: Normalized average penalty vs Number of Agents ($N$) where Number of channel is $M=1$ with success probability $0.95$.
@@ VII Numerical Results
  • Randomized Policy: This policy randomly selects $M$ agents. To select $M$ agents out of $N$ agents, we use a built-in MATLAB function called “randperm($N,M$)”.
  • Periodic Updating: The agents generate updates at every time slot and store in a FIFO queue. In our simulation, we consider the queue of each agent can store $1000$ updates. When the queue buffer is full, the oldest packet is dropped. At every time slot, $M$ agents are selected for transmission using the “randperm($N,M$)” function. Whenever an agent is selected, the oldest packet from the queue of the agent is sent to the receiver.
@@
We consider a safety-critical system with $N$ agents navigating within a region. This region is uniformly divided into a $20$ grid (20 rows and 20 columns). The observation $X_{n,t}$ for agent $n$ at time $t$ is represented by its position $x_{n}=(x_{n,1},x_{n,2})$. The safety level $Y_{n,t}$ is categorized into three regions: safe, cautious, and dangerous. Rows 1 through 6 are designated as safe, rows 7 through 13 as cautious, and the remaining rows as dangerous. This setup differs from the one depicted in Figure 4, where all $400$ positions represent distinct states, leading to a state space of size $\tau\times 400$, where $\tau$ is the truncated AoI value. This large state space significantly increases the complexity of determining the value function and gain index. To mitigate this, we simplify the representation by considering only the row information for safety assessment. Consequently, the total number of states is reduced from $\tau\times 400$ to $\tau\times 20$.
@@
Figure 6: Normalized average penalty vs Number of channels ($M$) where Number of agents are $N=20$ with success probability $0.95$.
@@
In this simulation, each agent $n$ can move randomly in any of the four directions: up, down, left, and right. We consider half of the agents move in up with probability $0.3$, down with probability $0.3$, left with probability $0.2$, and right with probability $0.2$. The rest of the agents move in up with probability $0.05$, down with probability $0.05$, left with probability $0.45$, and right with probability $0.45$. If an agent reaches a boundary, it stays in its current position. The loss incurred by agent $n$ is given by $L_{n}(y,\hat{y})$ which is defined as follows: $L_{n}$(dangerous, safe) $=1000,L_{n}$(safe, dangerous) $=5,L_{n}$(cautious, safe) $=10,L_{n}$(safe, cautious) $=1,1,L_{n}$(dangerous, cautious) $=100$, $L_{n}$(cautious, dangerous)$=5$, and $L_{n}$(dangerous, dangerous) $=L_{n}$(cautious, cautious) $=L_{n}$(safe, safe) $=0$. In our simulation, the success probability is $0.95$.
Figure 7: Normalized average penalty vs the scaling parameter ($r$) where the number of agents is $4r$ and the number of channels is $r$ with success probability $0.95$.
@@
Figure 5 illustrates the performance comparison of the four policies mentioned above as the number of agents increases. We consider two erasure channels in this simulation. The normalized average penalty (time-averaged penalty per agent) in Figure 5 is obtained by dividing time-average cost by the number of agents. From the figure, the MGF policy outperforms periodic updating, randomized policy, and MAF policy. The performance of periodic updating is always worse than the other three policies. This is because periodic updating sends packets even when the previous packets has not yet been delivered, resulting in an extended waiting time in the queue and a significantly large AoI value at the receiver side. Furthermore, the randomized policy randomly selects two agents for sending updates and the MAF policy only utilizes the AoI in decision-making. In contrast, the MGF policy makes the decision in a smarter way by considering the gain index, which is a function of both the AoI and latest received observation. Hence, the MGF policy shows better performance than the other three policies. The performance gain of the MGF policy is up to $10.47$ times compared to periodic updating, up to $2.33$ times compared to randomized policy, and up to $1.84$ times compared to MAF policy.
@@
Figure 6 illustrates the performance comparison of the four policies as the number of channels increases. We consider $20$ agents in this simulation. With the increase of the number of available channels for sending updates, the performance of the policies are getting better. However, because of the intelligent decision strategy by utilizing the AoI and the latest received observation, the MGF policy outperforms the other three policies. The performance gain of the MGF policy is up to $9.08$ times compared to periodic updating, up to $2.5$ times compared to randomized policy, and up to $1.96$ times compared to MAF policy.
@@
Figure 8: Penalty $q_{n}(\delta,x)$ vs received observation $x$ and Gain $\alpha_{n}(\delta,x)$ vs received observation $x$ for fixed AoI $\delta$. We consider number of agents $N=20$, number of channels $M=10$, success probability $0.95$, and $w_{n}=1$ for all agents
@@
Appendix A Proof of Theorem 1
@@
$\mathbb{P}(Y_{n,t}|\mathbf{H}_{n,t})$
$=$ $\mathbb{P}(Y_{n,t}|\big{(}X_{n,\tau-\Delta_{n}(\tau)},\Delta_{n}(\tau),\mu_{n}(\tau-1),\gamma_{n}(\tau-1)\big{)}_{\tau=0}^{t})$
@@
We will prove (A) in two steps: (1) First, we show that $Y_{n,t}$ is conditionally independent of the scheduling decisions $(\mu_{n}(\tau))_{\tau=0}^{t-1}$ and the delivery indicators $(\gamma_{n}(\tau))_{\tau=0}^{t-1}$ given the AoI history $(\Delta_{n}(\tau))_{\tau=0}^{t}$; (2) Next, we show that $Y_{n,t}$ is conditionally independent of the $(\Delta_{n}(\tau),X_{n,\tau-\Delta_{n}(\tau)})_{\tau=0}^{t-1}$ given the AoI $(\Delta_{n}(t),X_{n,t-\Delta_{n}(t)})$.
@@
Due to the Markov property of the underlying process $Y_{n,t}\leftrightarrow X_{n,t}\leftrightarrow X_{n,t-1}\leftrightarrow\ldots$, we have the following Markov chain given $\Delta_{n}(t)$: $Y_{n,t}\leftrightarrow X_{n,t-\Delta_{n}(t)}\leftrightarrow X_{n,(t-1)-\Delta_{n}(t-1)}\leftrightarrow\ldots$. Hence, $Y_{n,t}$ is independent of $(\Delta_{n}(\tau),X_{n,\tau-\Delta_{n}(\tau)})_{\tau=0}^{t-1}$ given $\Delta_{n}(t)$ and $X_{n,t-\Delta_{n}(t)}$. Considering these facts, (A) follows from (39). Therefore, we can write
$\inf_{y\in\mathcal{Y}_{n}}\mathbb{E}[L(Y_{n,t},y)|\mathbf{H}_{n,t}]$
$=$ $\inf_{y\in\mathcal{Y}_{n}}\!\!\!\mathbb{E}[L(Y_{n,t},y)|\Delta_{n}(t),X_{n,t-\Delta_{n}(t)}].$ (40)
@@
$=$ $\sum_{t=0}^{\infty}\sum_{n=1}^{N}\!\mathbb{E}\!\bigg{[}L_{n}(Y_{n,t},f_{n}(\Delta_{n}(t),X_{n,t-\Delta_{n}(t)})-\mathsf{L}_{\text{opt}}\bigg{|}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\Delta_{n}(t),X_{n,t-\Delta_{n}(t)}\bigg{]},$ (43)
where (A) holds from (A). Therefore, from [17, Chapter 4.3], Theorem 1 (ii) follows.
Appendix B Proof of Lemma 1
@@
$\sum_{(\delta,x)}$ $P_{\Delta_{n}(t),X_{n,t-\Delta_{n}(t)}}(\delta,x)\times$
$\inf_{f_{n}(x,\delta)\in\mathcal{Y}_{n}}\!\!\!\!\!\mathbb{E}_{P_{Y_{n,t}}|\Delta_{n}(t)=\delta,X_{n,t-\delta}=x}[L_{n}(Y,f_{n}(\delta,x)]$ (47)
$=\sum_{(\delta,x)}$ $P_{\Delta_{n}(t),X_{n,t-\Delta_{n}(t)}}(\delta,x)\times$
$\inf_{y_{n}\in\mathcal{Y}_{n}}\mathbb{E}_{P_{Y_{n,t}}|\Delta_{n}(t)=\delta,X_{n,t-\delta}=x}[L_{n}(Y,y_{n})].$ (48)
@@
Appendix C Proof of Lemma 2
@@
$H_{L}(Y_{n,t}|\Delta_{n}(t),X_{n,t-\Delta_{n}(t)})$
$=$ $\sum_{(\delta,x)\in\mathbb{Z}\times\mathcal{X}_{n}}P_{\Delta_{n}(t),X_{n,t-\Delta_{n}(t)}}(\delta,x)\times$
$\inf_{f_{n}(\delta,x)\in\mathcal{Y}_{n}}\mathbb{E}_{Y\sim P_{Y_{n,t}|\Delta_{n}(t)=\delta,X_{n,t-\delta}=x}}\big{[}L_{n}(Y,f_{n}(\delta,x))\big{]}$
$=$ $\sum_{(\delta,x)\in\mathbb{Z}\times\mathcal{X}_{n}}P_{\Delta_{n}(t),X_{n,t-\delta}}(\delta,x)\times$
$\inf_{f_{n}(\delta,x)\in\mathcal{Y}_{n}}\mathbb{E}_{Y\sim P_{Y_{n,t}|\Delta_{n}(t)=\delta,X_{n,t-\delta}=x}}\big{[}L_{n}(Y,f_{n}(\delta,x))\big{]}.$ (49)
Appendix D Proof of Theorem 2
@@
Let $\bar{u}^{n,\mu}_{\delta,x}$ be the optimal state action frequency of the Lagrangian problem (V-A) in which $\lambda=\lambda^{*}$ is the optimal dual Lagrangian multiplier.
@@
$\mathcal{S}_{-}^{n}=\{(\delta,x):\bar{u}_{\delta,x}^{n,1}<0{\thinspace}~{}{\text{and}}~{}{\thinspace}\bar{u}_{\delta,x}^{n,0}>0\}.$ (53)
@@
i) A class-$k$ bandit in state $(\delta_{k},x_{k})\in\mathcal{S}_{+}^{k}$ is given high priority than a class-$j$ bandit in state $(\delta_{j},x_{j})\in\mathcal{S}_{0}^{j}$.
ii) A class-$k$ bandit in state $(\delta_{k},x_{k})\in\mathcal{S}_{0}^{k}$ is given high priority than a class-$j$ bandit in state $(\delta_{j},x_{j})\in\mathcal{S}_{-}^{j}$.
iii) If a class-$k$ bandit is activated, then $\mu=1$ is chosen; otherwise, $\mu=0$ is chosen.
@@
By utilizing [57, Theorem 13], we can obtain Lemma 3. Following [57, Proposition 14] , for class-$k$ bandits
1. For any class-$k$ bandit, $(\delta_{k},x_{k})\in\mathcal{S}_{+}$ implies
@@ 3. For any class-$k$ bandit, $(\delta_{k},x_{k})\in\mathcal{S}_{-}$ implies
Because the policy $\pi_{\text{gain}}$ activates exactly $M$ bandits with the highest gain and if a bandit $k$ is activated, $\pi_{\text{gain}}$ chooses $\mu=1$, we can deduce form (54)-(56) and Definition 4 that $\pi_{\text{gain}}$ belongs to $\Pi_{\text{LP-Priority}}$. This concludes the proof.
Appendix E Special Case: Single-Source, Single-Channel

### v2

@@ V-C Asymptotic Optimality
$\Psi_{\pi}((\mathbf{v}^{n})_{n=1}^{N})=\mathbb{E}[(\mathbf{V}^{n}(t+1))_{n=1}^{N}|(\mathbf{V}^{n}(t))_{n=1}^{N}=(\mathbf{v}^{n})_{n=1}^{N}].$ (35)
@@ VI Discussions
[…]
Figure 5: Normalized average penalty vs Number of Agents ($N$) where Number of channel is $M=1$ with success probability 0.950.950.95.
@@ VII Numerical Results
[…]
    The difference between Randomized policy and Random Selection with Queue is: In Randomized policy, whenever an agent is selected for transmission, a fresh update with AoI=0 is sent; In Random Selection with Queue, updates are generated at every time-slots and stored in a queue. When an agent is selected for transmission, the oldest packet from the queue is sent.
@@
We consider a safety-critical system with $N$ agents navigating within a region. This region is uniformly divided into a 202020 grid (20 rows and 20 columns). The observation $X_{n,t}$ for agent $n$ at time $t$ is represented by its position $x_{n}=(x_{n,1},x_{n,2})$. The safety level $Y_{n,t}$ is categorized into three regions: safe, cautious, and dangerous. Rows 1 through 6 are designated as safe, rows 7 through 13 as cautious, and the remaining rows as dangerous. This setup differs from the one depicted in Figure 4, where all 400400400 positions represent distinct states, leading to a state space of size $\tau\times 400$, where $\tau$ is the truncated AoI value. This large state space significantly increases the complexity of determining the value function and gain index. To mitigate this, we simplify the representation by considering only the row information for safety assessment. Consequently, the total number of states is reduced from $\tau\times 400$ to $\tau\times 20$.
@@
Figure 6: Normalized average penalty vs Number of channels ($M$) where Number of agents are $N=20$ with success probability 0.950.950.95.
@@
In this simulation, each agent $n$ can move randomly in any of the four directions: up, down, left, and right. We consider half of the agents move in up with probability 0.30.30.3, down with probability 0.30.30.3, left with probability 0.20.20.2, and right with probability 0.20.20.2. The rest of the agents move in up with probability 0.050.050.05, down with probability 0.050.050.05, left with probability 0.450.450.45, and right with probability 0.450.450.45. If an agent reaches a boundary, it stays in its current position. The loss incurred by agent $n$ is given by $L_{n}(y,\hat{y})$ which is defined as follows: $L_{n}$(dangerous, safe) $=1000,L_{n}$(safe, dangerous) $=5,L_{n}$(cautious, safe) $=10,L_{n}$(safe, cautious) $=1,1,L_{n}$(dangerous, cautious) =100=100= 100, $L_{n}$(cautious, dangerous)=5=5= 5, and $L_{n}$(dangerous, dangerous) $=L_{n}$(cautious, cautious) $=L_{n}$(safe, safe) =0=0= 0. In our simulation, the success probability is 0.950.950.95.
[…]
@@
Figure 5 illustrates the performance comparison of the four policies mentioned above as the number of agents increases. We consider two erasure channels in this simulation. The normalized average penalty (time-averaged penalty per agent) in Figure 5 is obtained by dividing time-average cost by the number of agents. From the figure, the MGF policy outperforms random selection with queue, randomized policy, and MAF policy. The performance of random selection with queue is always worse than the other three policies. This is because random selection with queue sends packets even when the previous packets has not yet been delivered, resulting in an extended waiting time in the queue and a significantly large AoI value at the receiver side. Furthermore, the randomized policy randomly selects two agents for sending updates and the MAF policy only utilizes the AoI in decision-making. In contrast, the MGF policy makes the decision in a smarter way by considering the gain index, which is a function of both the AoI and latest received observation. Hence, the MGF policy shows better performance than the other three policies. The performance gain of the MGF policy is up to 10.4710.4710.47 times compared to random selection with queue, up to 2.332.332.33 times compared to randomized policy, and up to 1.841.841.84 times compared to MAF policy.
@@
Figure 6 illustrates the performance comparison of the four policies as the number of channels increases. We consider 202020 agents in this simulation. With the increase of the number of available channels for sending updates, the performance of the policies are getting better. However, because of the intelligent decision strategy by utilizing the AoI and the latest received observation, the MGF policy outperforms the other three policies. The performance gain of the MGF policy is up to 9.089.089.08 times compared to random selection with queue, up to 2.52.52.5 times compared to randomized policy, and up to 1.961.961.96 times compared to MAF policy.
@@
Figure 8: Penalty $q_{n}(\delta,x)$ vs received observation $x$ and Gain $\alpha_{n}(\delta,x)$ vs received observation $x$ for fixed AoI $\delta$. We consider number of agents $N=20$, number of channels $M=10$, success probability 0.950.950.95, and $w_{n}=1$ for all agents.
@@
Appendix A Definitions of 0-1 loss, quadratic loss, logarithmic loss
0-1 loss: The 0-111 loss function assigns a loss of 0 for incorrect estimation of a random variable $Y=y$, and 111 otherwise. It is given by and is given by an indicator function which equals to 1 if the true value $y$ and the estimated value $\hat{y}$ are equal to each other. $L_{0-1}(y,\hat{y})$ is given by
$L_{0-1}(y,\hat{y})=\mathbf{1}\{y\neq\hat{y}\},$ (38)
where $\mathbf{1}\{y\neq\hat{y}\}$ is the indicator function for the event $\{y\neq\hat{y}\}$.
Quadratic loss: The quadratic loss function quantifies the error between the true value of a random variable $Y=y$ and and its estimated value $\hat{y}$ by calculating the square of their difference. It is given by
$L_{2}(y,\hat{y})=(y-\hat{y})^{2}.$ (39)
Logarithmic loss: The log-loss function $L_{\log}(y,P_{Y})$ is the negative log-likelihood of the true value $Y=y$ which is given by
$L_{\log}(y,P_{Y})=-\log P_{Y}(y)$ (40)
where the action $a=P_{Y}$ is a distribution of $Y$.
Appendix B Proof of Theorem 1
@@
$\mathbb{P}(Y_{n,t}|\mathbf{H}_{n,t})=$ $\mathbb{P}(Y_{n,t}|\big{(}X_{n,\tau-\Delta_{n}(\tau)},\Delta_{n}(\tau),\mu_{n}(\tau-1),\gamma_{n}(\tau-1)\big{)}_{\tau=0}^{t})$
@@
We will prove (B) in two steps: (1) First, we show that $Y_{n,t}$ is conditionally independent of the scheduling decisions $(\mu_{n}(\tau))_{\tau=0}^{t-1}$ and the delivery indicators $(\gamma_{n}(\tau))_{\tau=0}^{t-1}$ given the AoI history $(\Delta_{n}(\tau))_{\tau=0}^{t}$; (2) Next, we show that $Y_{n,t}$ is conditionally independent of the $(\Delta_{n}(\tau),X_{n,\tau-\Delta_{n}(\tau)})_{\tau=0}^{t-1}$ given the AoI $(\Delta_{n}(t),X_{n,t-\Delta_{n}(t)})$.
@@
Due to the Markov property of the underlying process $Y_{n,t}\leftrightarrow X_{n,t}\leftrightarrow X_{n,t-1}\leftrightarrow\ldots$, we have the following Markov chain given $\Delta_{n}(t)$: $Y_{n,t}\leftrightarrow X_{n,t-\Delta_{n}(t)}\leftrightarrow X_{n,(t-1)-\Delta_{n}(t-1)}\leftrightarrow\ldots$. Hence, $Y_{n,t}$ is independent of $(\Delta_{n}(\tau),X_{n,\tau-\Delta_{n}(\tau)})_{\tau=0}^{t-1}$ given $\Delta_{n}(t)$ and $X_{n,t-\Delta_{n}(t)}$. Considering these facts, (B) follows from (42). Therefore, we can write
$\inf_{y\in\mathcal{Y}_{n}}\mathbb{E}[L(Y_{n,t},y)|\mathbf{H}_{n,t}]=\inf_{y\in\mathcal{Y}_{n}}\!\!\!\mathbb{E}[L(Y_{n,t},y)|\Delta_{n}(t),X_{n,t-\Delta_{n}(t)}].$ (43)
@@
$=$ $\sum_{t=0}^{\infty}\sum_{n=1}^{N}\!\mathbb{E}\!\bigg{[}L_{n}(Y_{n,t},f_{n}(\Delta_{n}(t),X_{n,t-\Delta_{n}(t)})-\mathsf{L}_{\text{opt}}\bigg{|}\Delta_{n}(t),X_{n,t-\Delta_{n}(t)}\bigg{]},$ (46)
where (B) holds from (B). Therefore, from [17, Chapter 4.3], Theorem 1 (ii) follows.
Appendix C Proof of Lemma 1
@@
$\sum_{(\delta,x)}P_{\Delta_{n}(t),X_{n,t-\Delta_{n}(t)}}(\delta,x)\inf_{f_{n}(x,\delta)\in\mathcal{Y}_{n}}\!\!\!\!\!\mathbb{E}_{P_{Y_{n,t}}|\Delta_{n}(t)=\delta,X_{n,t-\delta}=x}[L_{n}(Y,f_{n}(\delta,x)]$ (50)
$=$ $\sum_{(\delta,x)}P_{\Delta_{n}(t),X_{n,t-\Delta_{n}(t)}}(\delta,x)\inf_{y_{n}\in\mathcal{Y}_{n}}\mathbb{E}_{P_{Y_{n,t}}|\Delta_{n}(t)=\delta,X_{n,t-\delta}=x}[L_{n}(Y,y_{n})].$ (51)
@@
Appendix D Proof of Lemma 2
@@
$H_{L}(Y_{n,t}|\Delta_{n}(t),X_{n,t-\Delta_{n}(t)})=$ $\sum_{(\delta,x)\in\mathbb{Z}\times\mathcal{X}_{n}}P_{\Delta_{n}(t),X_{n,t-\Delta_{n}(t)}}(\delta,x)\times\inf_{f_{n}(\delta,x)\in\mathcal{Y}_{n}}\mathbb{E}_{Y\sim P_{Y_{n,t}|\Delta_{n}(t)=\delta,X_{n,t-\delta}=x}}\big{[}L_{n}(Y,f_{n}(\delta,x))\big{]}$
$=$ $\sum_{(\delta,x)\in\mathbb{Z}\times\mathcal{X}_{n}}P_{\Delta_{n}(t),X_{n,t-\delta}}(\delta,x)\inf_{f_{n}(\delta,x)\in\mathcal{Y}_{n}}\mathbb{E}_{Y\sim P_{Y_{n,t}|\Delta_{n}(t)=\delta,X_{n,t-\delta}=x}}\big{[}L_{n}(Y,f_{n}(\delta,x))\big{]}.$ (52)
Appendix E Proof of Theorem 2
@@
Let $\bar{u}^{n,\mu}_{\delta,x}$ be the optimal state action frequency of the Lagrangian problem (25) in which $\lambda=\lambda^{*}$ is the optimal dual Lagrangian multiplier.
@@
$\mathcal{S}_{-}^{n}=\{(\delta,x):\bar{u}_{\delta,x}^{n,1}=0{\thinspace}~{\text{and}}~{\thinspace}\bar{u}_{\delta,x}^{n,0}\geq 0\}.$ (56)
@@
[…]
@@
By utilizing [57, Theorem 13], we can obtain Lemma 3. The results in [57] hold for RMAB problem with equality constraint. Our problem (4)-(5) considers an inequality constraint (5). In order to use the results from [57], we introduce dummy bandits and obtain an equivalent problem (16)-(19) with equality constraint (17).
[…]
@@
Because the policy $\pi_{\text{gain}}$ activates exactly $rM$ bandits with the highest gain and if a bandit $k$ is activated, $\pi_{\text{gain}}$ chooses $\mu=1$, we can deduce form (57)-(59) and Definition 4 that $\pi_{\text{gain}}$ belongs to $\Pi_{\text{LP-Priority}}$. This concludes the proof.
Appendix F Special Case: Single-Source, Single-Channel